OpenStudy (anonymous):

how do I solve this problem x=0.6x+384

OpenStudy (anonymous):

Easier to think of it as: 1x=0.6x+384 Combine your X's by subtracting 0.4x on both sides to get: 0.4x=384 Divide by 0.4 to get x alone on one side: x=960

OpenStudy (anonymous):

how did you get 0.4x

OpenStudy (anonymous):

I don't understand how you came up with 0.4x

OpenStudy (anonymous):

1x=0.6x+384 -.6x from both sides .4x = 384

OpenStudy (anonymous):

Yes, Like DingusExpress said :) 1x = 0.6x + 384 -0.6x -0.6x _________________ 0.4x = 0 + 384 __________________ Then you divide on both sides by the coefficient of x, which is 0.4, and you get x = 960